OpenStudy (anonymous):

Jason and Ron together have 97 baseball cards. Jason has four more than twice as many cards as ron. How many baseball cards does each person have?

OpenStudy (anonymous):

set it up as equations. jason has 4 MORE THAN TWICE as many as ron. we know more than means? and that twice means?

OpenStudy (anonymous):

Let's say Jason has "x" amount of cards and Ron has "y" amount of cards. Equation 1) Together, they have 97 cards, so say: x + y = 97 Equation 2) Jason has 4 more than twice as many cards as Ron. Jason's cards = (Twice Ron's cards) + 4 Twice Ron's cards is the same thing as "2y," so just put that and the symbol for Jason's cards into the equation --> x = (2y) + 4 So your two equations are: 1) x + y = 97 2) x = 2y + 4 One of the equations tell you what x is equal to, so just sub the second equation into the first. Equation 1 changes from: x + y = 97 --> (2y + 4) + y = 97 Combine like terms and simplify to give you: 3y + 4 = 97 Subtract 4 from both sides then divide both sides by 3: y = 31 So now we know Ron has 31. We also know both their cards combined is 97, so just subtract 31 from 97 to find how many cards Jason has: 97 - 31 = ?

OpenStudy (anonymous):

66